A 20 metre length of cable has a cross-sectional area of 1mm2and a resistance of 5 ohms.  Calculate the conductivity of the cable. 



1
Expert's answer
2021-08-30T15:11:41-0400

Gives

l=20m

Area(A)=1mm2

R=5ohm

We know that

"R=\frac{\rho l}{A}"

"\rho=\frac{RA}{l}"

Put value


"\rho=\frac{5\times10^{-6}}{20}=2.5\times10^{-7}\Omega m"

We know that

"\sigma=\frac{1}{\rho}"


"\sigma=\frac{1}{2.5\times10^{-7}}=4\times10^6mho/m"